# Perverse microsheaves

### v2

@@ 2. Complex contact and symplectic manifolds
Observe that there is a fiber-preserving $\mathbb{C}^{*}$-action on $\tilde{V}$: over some fiber $\tilde{V}_{v}\subset\Omega_{V,v}$, it sends $\alpha\mapsto z\alpha$ for $z\in\mathbb{C}^{*}$ and $v\in V$. For concreteness, we write $z=e^{t+i\theta}$ for $(t,\theta)\in\mathbb{R}_{+}\times S^{1}$ and let $\partial_{t},\partial_{\theta}$ denote the vector fields on $\tilde{V}$ generated by the $\mathbb{R}_{+}$ and $S^{1}$ actions.
We have the following morphisms of bundles over $\tilde{V}$:
(7) $\pi^{*}\xi\equiv\operatorname{ker}(\lambda_{\tilde{V}})/\langle\partial_{t},\partial_{\theta}\rangle\twoheadleftarrow\operatorname{ker}(\lambda_{\tilde{V}})\hookrightarrow\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))\hookrightarrow T\tilde{V}$
@@
There is a splitting of vector bundles over $\tilde{V}$ (well-defined up to contractible choice)
$\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))=\partial_{t}\oplus\langle\partial_{\theta},X\rangle\oplus\pi^{*}\xi,$
where $X$ is a non-vanishing section of $\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\tilde{V}}))/\operatorname{ker}(\hbar\lambda_{\tilde{V}})$ and
@@
    $\langle\partial_{\theta},X\rangle$ is a trivial $2$-dimensional real symplectic vector bundle with respect to $d(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$
@@
    $\pi^{*}\xi$ is a complex symplectic vector bundle with respect to $d(\lambda_{\tilde{V}})$
@@
Given a complex contact manifold $(V,\mathcal{H})$ of complex dimension $12n+$ 1, a complex submanifold $L\subset V$ of complex dimension $n$ which is everywhere tangent to $\mathcal{H}$ is said to be a complex Legendrian.
@@
If $L\subset(V,\xi)$ is complex Legendrian, then its preimage under $\tilde{V}\to V$ is denoted by $\tilde{L}$ and is complex exact Lagrangian. The quotient $\tilde{L}/\mathbb{R}_{+}\subset\tilde{V}/\mathbb{R}_{+}$ is a real Legendrian with respect to any of the $\mathbb{C}^{*}$ of contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
@@ 3. Grading and orientation data
The material in this section is mostly algebraic topology. We will need it in Section 5 when we discuss how the condition of being a complex contact manifold/complex Legendrian interacts with orientations/gradings in microlocal sheaf theory.
@@
Here $\sqrt{SU}(n)$ is defined as the kernel of $U(n)\xrightarrow{det^{2}}U(1)$. Note that $U(n)$ is the maximal compact subgroup of both $Sp(2n,\mathbb{R})$ and $GL(n,\mathbb{C})$. Meanwhile $U(n/2,\mathbb{H})$ is also known as the ‘compact symplectic group’ $Sp(n/2)=Sp(n,\mathbb{C})\cap U(n)$, the maximal compact subgroup of $Sp(n,\mathbb{C})$.
@@
  • A grading (or grading datum) is a lift to $B(\sqrt{SU}(n))$
@@
Remark 3.3.
We emphasize that the identifications in Definition 3.1 should be understood as holding up to coherent homotopy. So, for example, a polarization for $X\to BU(n)$ is a map $X\to BO(n)$ and a homotopy between the composition $X\to BO(n)\to BU(n)$ and the original structure map $X\to BU(n)$.
Remark 3.4.
Strictly speaking, a symplectic vector bundle $E$ on $X$ is classified by a map $X\to BSp(2n,\mathbb{R})$. However, $BU(n)\to BSp(2n,\mathbb{R})$ is a homotopy equivalence and we may always lift to some $X\to BU(n)$ uniquely up to contractible choice. (I.e. the space of Hermitian structures compatible with a given symplectic bundle is contractible.) Any Lagrangian distribution in $E$ will be totally real for any compatible Hermitian structure, hence define a reduction of structure to $O(n)$, i.e. what we have here called a polarization.
Likewise, a complex symplectic bundle carries a quaternionic structure, uniquely up to contractible choices. We typically omit discussion of such issues in the sequel. In particular, for $(X,\omega)$ is complex symplectic, we simply say $TX$ carries a quaternionic structure.
Remark 3.5.
@@
The sequence $O(n)\to U(n)\xrightarrow{det^{2}}S^{1}$ is a fiber sequence, which we may deloop to a fiber sequence $BO(n)\to BU(n)\xrightarrow{Bdet^{2}}BS^{1}$. So, a symplectic vector bundle $E$ admits a grading iff $02c_{1}(E)=$ 0. In particular, a symplectic manifold $X$ admits a grading iff $02c_{1}(X)=$ 0.
@@
$Sp=\lim_{n\to\infty}U(n,\mathbb{H})\qquad\qquad U=\lim_{n\to\infty}U(n,\mathbb{C})\qquad\qquad O=\lim_{n\to\infty}O(n,\mathbb{R})$
and the Lagrangian Grassmannians
@@
There are evident stable analogues of the notions of Def. 3.1, and the stable analogue of Lemma 3.2 also holds.
@@
Let $(V,\xi)$ be a complex contact manifold and consider the real contact manifold $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. Then the contact distribution $\xi_{\hbar}\to\tilde{V}/\mathbb{R}_{+}$ carries a stable quaternionic structure.
@@
It is equivalent to prove that $q^{*}\xi_{\hbar}$ carries a stable quaternionic structure, where $q:\tilde{V}\to\tilde{V}/\mathbb{R}_{+}$ is the quotient map. This follows from Lemma 2.7. ∎
We write $w_{2}:BO\to B^{2}(\mathbb{Z}/2\mathbb{Z})$ for the map classified by the second Stiefel-Whitney class.
@@
Let $X$ be a topological space and consider a map $\nu:X\to BU$, i.e. a stable complex or symplectic vector bundle on $X$. Orientation data is a null-homotopy of the composite map:
@@
The canonical orientation data is the null-homotopy arising from the canonical null-homotopy of $BU\to B(U/O)\to B^{2}O$.
Evidently, the space of orientation data is identified (using the canonical orientation datum) with $Maps(X,B^{2}(\mathbb{Z}/2\mathbb{Z}))$; in particular, the choices up to homotopy are $H^{2}(X,\mathbb{Z}/2\mathbb{Z})$.
[…]
(8) $X\to BO\to BU\to B(U/O)\to B^{2}O\xrightarrow{Bw_{2}}B^{3}(\mathbb{Z}/2\mathbb{Z})$
there are two evident null-homotopies, coming from the compositions $BO\to BU\to B(U/O)$ and $BU\to B(U/O)\to B^{2}O$. We term the first the polarization orientation data; the second is what we called above the canonical orientation data.
@@
Fix a stable polarization $X\to BO$ of a given $X\to BU$. The space of homotopies between the canonical and polarization orientation data is equivalent to the space of lifts to $X\to BPin$.
Proof.
First we recall some general facts about null-homotopies and exact triangles. Given two null-homotopies $n_{1},n_{2}$ of a given map $f:Q\to R$, we produce a pointed map $[n_{1},n_{2}]\in Q\times S^{1}\to R$ by taking $S^{1}=[-\pi,\pi]$, taking the map $f$ on $Q\times 0$ and applying the null-homotopy $n_{1}$ along $[-\pi,0]$ and the null-homotopy $n_{2}$ along $[0,\pi]$. Note that $\operatorname{Hom}(Q\times S^{1},R)=\operatorname{Hom}(Q,\Omega R)$. A homotopy between $n_{1}$ and $n_{2}$ is a null-homotopy of $[n_{1},n_{2}]\in\operatorname{Hom}(Q,\Omega R)$.
Suppose now given any exact triangle $P\to Q\to R\to BP$ in a stable category (for the application here, the stable category of spectra). The compositions $P\to Q\to R$ and $Q\to R\to BP$ give two null-homotopies of the composite map $P\to BP$. The definition of exact triangles [19, Def. 1.1.2.11] promises that the comparison of these two null-homotopies is identified with the identity of $\operatorname{Hom}(P,\Omega BP)=\operatorname{Hom}(P,P)$.
[…]
given by $[s\circ n_{1}\circ p,s\circ n_{2}\circ p]\in\operatorname{Hom}(X,\Omega BS)$ is identified with $s\circ p\in\operatorname{Hom}(X,S)$. Here, $X$ need only be a space, not a spectrum.
Now specializing to (8), we find that the space of homotopies between polarization and canonical orientation data is identified with the space of null-homotopies of $X\to BO\xrightarrow{w_{2}}B^{2}(\mathbb{Z}/2\mathbb{Z})$, i.e., the space of lifts to $X\to BPin$. ∎
Definition 3.10.
Let $V$ be a symplectic or contact manifold. Fix a polarization / grading / orientation data $m$ on $V$.
Consider $L\subset V$ a Lagrangian or Legendrian. There is a canonical isomorphism of $T^{*}L$ with the restriction of the symplectic tangent bundle or contact distribution. We write $\phi_{L}$ for the fiber polarization.
A relative polarization for $L$ is a homotopy between $m|_{L}$ and $\phi_{L}$. Relative grading / orientation data are similarly homotopies between the grading / orientation data induced by $m|_{L}$ and $\phi_{L}$.
Note that the obstruction to existence of a relative grading is a class in $[L,B\mathbb{Z}]=H^{1}(L,\mathbb{Z})$, and the space of relative gradings is a torsor for $Map(L,\mathbb{Z})$, hence in particular, is discrete. Thus we simply ask whether relative gradings are equal, rather than discuss homotopies between them.
Lemma 3.11.
Let $V$ is a complex contact manifold, $\tilde{V}\to V$ the corresponding real contact manifold equipped with the canonical grading constructed in Lemma 3.2. Let $L\subset V$ be a smooth complex Legendrian. Then there is a canonical choice for the relative grading of $\tilde{L}\subset\tilde{V}$.
@@
Fix the canonical orientation data on $V$. Then relative orientation data for $L\subset V$ is equivalent to a pin structure on $L$. (If $L$ is oriented, then a pin structure on $L$ is equivalent to a spin structure on $L$ compatible with the given orientation.)

### v3

@@ 2. Complex contact and symplectic manifolds
Observe that there is a fiber-preserving $\mathbb{C}^{*}$-action on $\widetilde{V}$: over some fiber $\widetilde{V}_{v}\subset\Omega_{V,v}$, it sends $\alpha\mapsto z\alpha$ for $z\in\mathbb{C}^{*}$ and $v\in V$. For concreteness, we write $z=e^{t+i\theta}$ for $(t,\theta)\in\mathbb{R}_{+}\times S^{1}$ and let $\partial_{t},\partial_{\theta}$ denote the vector fields on $\widetilde{V}$ generated by the $\mathbb{R}_{+}$ and $S^{1}$ actions.
We have the following morphisms of bundles over $\widetilde{V}$:
(7) $\pi^{*}\xi\equiv\operatorname{ker}(\lambda_{\widetilde{V}})/\langle\partial_{t},\partial_{\theta}\rangle\twoheadleftarrow\operatorname{ker}(\lambda_{\widetilde{V}})\hookrightarrow\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\widetilde{V}}))\hookrightarrow T\widetilde{V}$
@@
There is a splitting of vector bundles over $\widetilde{V}$ (well-defined up to contractible choice)
$\operatorname{ker}(\operatorname{re}(\hbar\lambda_{\widetilde{V}}))=\partial_{t}\oplus\langle\partial_{\theta},X\rangle\oplus\pi^{*}\xi,$
[…]
@@
    $\langle\partial_{\theta},X\rangle$ is a trivial $2$-dimensional real symplectic vector bundle with respect to $)d(\operatorname{re}(\hbar\lambda_{\widetilde{V}}$))
@@
[…]
@@
Given a complex contact manifold $(V,\mathcal{H})$ of complex dimension $2n+1$, a complex submanifold $L\subset V$ of complex dimension $n$ which is everywhere tangent to $\mathcal{H}$ is said to be a complex Legendrian.
@@
If $L\subset(V,\xi)$ is complex Legendrian, then its preimage under $\widetilde{V}\to V$ is denoted by $\widetilde{L}$ and is complex exact Lagrangian. The quotient $\widetilde{L}/\mathbb{R}_{+}\subset\widetilde{V}$$\mathbb{R}_{+}$ is a real Legendrian with respect to any of the $\mathbb{C}^{*}$ of contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.
@@ 3. Grading and orientation data
Here $\sqrt{SU}(n)$ is defined as the kernel of $U(n)\xrightarrow{det^{2}}U(1)$. Recall that $U(n)$ is the maximal compact subgroup of both $Sp(2n,\mathbb{R})$ and $GL(n,\mathbb{C})$, which retract to it; correspondingly, $BU(n)$ is canonically homotopy equivalent to the spaces classifying either complex or symplectic vector bundles. Meanwhile $U(n/2,\mathbb{H})$ is also known as the ‘compact symplectic group’ $Sp(n/2)=Sp(n,\mathbb{C})\cap U(n)$, the maximal compact subgroup of $Sp(n,\mathbb{C})$.
@@
  • A grading is a lift to $B(\sqrt{SU}(n))$
@@
Concretely, a polarization of a symplectic vector bundle can be viewed as Lagrangian plane field, as follows by inspecting the fiber sequence $O(n)\to U(n)\to U(n)/O(n)=LGr(n)$. Similarly, a complex polarization of a complex symplectic vector bundle can be viewed as a complex Lagrangian plane field. We freely pass between both viewpoints in this paper.
@@
We will refer to gradings induced by any of the above structures as canonical gradings.
@@
A grading is a lift from $X\to BU(n)$ to $X\to B\sqrt{SU}(n)$, so equivalently, a null-homotopy of the composition $X\to BU(n)\xrightarrow{Bdet^{2}}BU(1)$. Recalling that homotopy class of $Bdet:BU(n)\to BU(1)=B^{2}\mathbb{Z}$ in $BU(n),B^{2}\mathbb{Z}]=\mathrm{H}^{2}(BU(n),\mathbb{Z})$ is the universal first Chern class, we see that the obstruction to the existence of a grading is twice the first Chern class.
@@
$Sp=\lim_{n\to\infty}U(n,\mathbb{H})\qquad\qquad U=\lim_{n\to\infty}U(n,\mathbb{C})\qquad\qquad O=\lim_{n\to\infty}O(n,\mathbb{R}).$
We write similarly $SU,\sqrt{SU}$. Note the inclusion $U(n,\mathbb{C})\to O(2n,\mathbb{R})$ lands in $SO(2n,\mathbb{R})$; correspondingly we have inclusions $U\subset SO\subset O$. We also have the Lagrangian Grassmannians
@@
There are evident stable analogues of the notions of Definition 3.1, and the stable analogue of Lemma 3.2 also holds.
@@
Let $(V,\xi)$ be a complex contact manifold and consider the real contact manifold $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. Then the contact distribution $\xi_{\hbar}\to\widetilde{V}$$\mathbb{R}_{+}$ carries a stable quaternionic structure.
@@
It is equivalent to prove that $q^{*}\xi_{\hbar}$ carries a stable quaternionic structure, where $q:\widetilde{V}\to\widetilde{V}$$\mathbb{R}_{+}$ is the quotient map. This follows from Lemma 2.7. ∎
One virtue of having passed to stabilizations is that $U/O$ is an infinite loop space. The map $det^{2}$ descends to an equivalence $det^{2}:\tau_{\leq 1}(U/O)\xrightarrow{\sim}B\mathbb{Z}$. Thus a grading on $X\to BU(n)$ is equivalently a null-homotopy of the composition $X\to BU(n)\to BU\to B(U/O)\to\tau_{\leq 2}B(U/O)$.
@@
Grading/orientation data for $X\to BU(n)$ is a null-homotopy of the composition $X\to BU(n)\to\tau_{\leq 3}B(U/O)$.
Note that a stable polarization for $X\to BU(n)$ is a null-homotopy of the composition $X\to BU\to B(U/O)$, so canonically provides grading/orientation data. We refer to such grading/orientation data as polarization grading/orientation data.
Truncations give the fiber sequence (of infinite loop spaces)
$B^{3}(\mathbb{Z}/2\mathbb{Z})\to\tau_{\leq 3}B(U/O)\to\tau_{\leq 2}B(U/O)=B\left(\tau_{\leq 1}(U/O)\right)\stackrel{{\scriptstyle Bdet^{2}}}{{=}}B^{2}\mathbb{Z}.$
We refer to the space of grading/orientation data lifting a given grading as orientation data.
Remark 3.6.
It is presumably implicit in the construction of orientations on moduli spaces for Floer theory that there is a universal way to choose orientation data on all (real) symplectic manifolds. In fact there are two such ways; and their existence is somewhat subtle from our present viewpoint, as pointed out to us by Sanath Devalapurkar. The main source of the subtlety is the fact that the fiber sequence $B(\mathbb{Z}/2\mathbb{Z})\to\pi_{\leq 1}(U/O)\to\mathbb{Z}$ splits as a sequence of topological spaces, but not as a sequence of infinite loop spaces; in fact, $B^{3}(\mathbb{Z}/2\mathbb{Z})\to B^{2}\pi_{\leq 1}(U/O)\to B^{2}(\mathbb{Z})$ does not split as a sequence of spaces. We record his arguments in Appendix LABEL:sanath. This fact is not logically required for the present article, because the existence of orientation data for complex symplectic manifolds is less subtle, as will be shown presently.
Lemma 3.7.
The composition $Sp\to U\to U/O\to\tau_{\leq 2}(U/O)$ factors canonically through $\tau_{\leq 2}(Sp)=0$.
Definition 3.8.
Given a stable quaternionic bundle $X\to BSp$, the canonical grading/orientation datum of the induced complex bundle $X\to BSp\to BU$ is the null-homotopy induced from the null-homotopy of $Sp\to\tau_{\leq 2}(U/O)$ above. Similarly, the canonical grading is the null-homotopy induced from the similar null-homotopy of $Sp\rightarrow\tau_{\leq 1}(U/O)$.
We note that the grading induced by the canonical grading/orientation datum agrees with the canonical grading of Lemma 3.2.
Because $Sp\subset\sqrt{SU}$, the map $Sp/U\to U/O\xrightarrow{det^{2}}U(1)=\tau_{\leq 1}(U/O)$ is canonically null-homotopic. So the map $Sp/U\to U/O\to\tau_{\leq 2}(U/O)$ lifts to $Sp/U\to B^{2}(\mathbb{Z}/2\mathbb{Z})$.
Lemma 3.9.
This lift is the composition $Sp/U\to BU\xrightarrow{c_{1}}B^{2}\mathbb{Z}\to B^{2}(\mathbb{Z}/2\mathbb{Z})$, where $c_{1}$ is the first Chern class.
Proof.
Since $\tau_{\leq 2}Sp=0$, any map $Sp/U\to B^{2}(\mathbb{Z}/2\mathbb{Z})$ must factor through $BU$. There’s a unique nontrivial map in $BU,B^{2}(\mathbb{Z}/2\mathbb{Z})]$, and it’s the reduction mod 2 of the first Chern class. Finally, it is standard that $\tau_{\leq 2}(Sp/U\to U/O)$ is nontrivial: one has the map between fiber sequences
@@
which induces a map between the long exact sequence of homotopy groups, which shows that $\pi_{2}(Sp/U)=\pi_{1}(U)=\mathbb{Z}$, $\pi_{2}(U/O)=\pi_{1}(O)=\mathbb{Z}/2$, and the map $\pi_{2}(Sp/U)\rightarrow\pi_{2}(U/O)$ is given by the quotient map $\mathbb{Z}\rightarrow\mathbb{Z}/2$. ∎
Proposition 3.10.
Fix a stable complex polarization $X\to BU$ of a given $X\to BSp$. The space of homotopies between the canonical and polarization orientation data is equivalent to the space of null-homotopies of the composition $X\to BU\xrightarrow{c_{1}}B^{2}\mathbb{Z}\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$.
Proof.
First we recall some general facts about null-homotopies and exact triangles. Given two null-homotopies $n_{1},n_{2}$ of a given map $f:Q\to R$, we produce a pointed “comparison map” $n_{1},n_{2}]\in Q\times S^{1}\to R$ by taking $S^{1}=[-\pi,\pi]$, taking the map $f$ on $Q\times 0$ and applying the null-homotopy $n_{1}$ along $\pi,0]$ and the null-homotopy $n_{2}$ along $0,\pi]$. Note that $\operatorname{Hom}(Q\times S^{1},R)=\operatorname{Hom}(Q,\Omega R)$. A homotopy between $n_{1}$ and $n_{2}$ is a null-homotopy of $n_{1},n_{2}]\in\operatorname{Hom}(Q,\Omega R)$.
Suppose now given any exact triangle $P\to Q\to R\to BP$ in a stable category (for the application here, the stable category of spectra). The compositions $P\to Q\to R$ and $Q\to R\to BP$ give two null-homotopies of the composite map $P\to BP$. The definition of exact triangles [30, Def. 1.1.2.11] promises that the comparison of these two null-homotopies is identified with the identity of $\operatorname{Hom}(P,\Omega BP)=\operatorname{Hom}(P,P)$.
Given maps $X\xrightarrow{p}P\xrightarrow{s}S$, we may compose to learn that the corresponding comparison between null-homotopies of
$X\xrightarrow{p}P\to Q\to R\to BP\xrightarrow{Bs}BS$
given by $s\circ n_{1}\circ p,s\circ n_{2}\circ p]\in\operatorname{Hom}(X,\Omega BS)$ is identified with $s\circ p\in\operatorname{Hom}(X,S)$. Here, $X$ need only be a space, not a spectrum.
Now consider a stable quaternionic vector bundle with a stable complex polarization, i.e. we have a lift $X\to BU\to BSp$. Consider the composition
(8) $X\to BU\to BSp\to B(Sp/U)\to B^{2}U\xrightarrow{Bc_{1}}B^{3}\mathbb{Z}\rightarrow B^{3}(\mathbb{Z}/2\mathbb{Z})$
Now, the canonical orientation data factors through $BSp\to 0$, hence is induced by the null-homotopy of the sequence $BSp\to B(Sp/U)\to B^{2}U$. The polarization orientation data comes from the prescribed null-homotopy of $X\to B(U/O)$; since we have a complex polarization this factors through a null-homotopy of $X\to B(Sp/U)$, hence is induced from the null-homotopy of $BU\to BSp\to B(Sp/U)$. The result follows. ∎
We recall that the mod 2 reduction of the first Chern class of a complex bundle is the second Stiefel-Whitney class $w_{2}$ of the underlying real bundle. Thus, null-homotopies of $X\to BU\xrightarrow{w_{2}}B^{2}(\mathbb{Z}/2\mathbb{Z})$ as above are the same as spin structures on the underlying real bundle classified by $X\to BU\to BSO$.
Let $V$ be a symplectic (resp. contact) manifold. Given a Lagrangian (resp. Legendrian) submanifold, the Weinstein neighborhood theorem symplectomorphically (resp. contactomorphically) identifies a neighborhood of $L\subset V$ with a neighborhood of the zero section in $T^{*}L$ (resp. in $J^{1}L$). This identification is canonical up to contractible choice. We write $\phi_{L}$ for the fiber polarization of $T^{*}L$ (resp. $J^{1}L$).
Definition 3.11.
Let $m$ be a grading (resp. grading/orientation datum) on $V$. Then by a secondary grading (resp. grading/orientation datum) for $L\subset V$, we mean a homotopy between $m|_{L}$ and $\phi_{L}$.
Given grading/orientation data on $V$ and a secondary grading on $L$, then by secondary orientation data, we mean a lift of said secondary grading on $L$ to secondary grading/orientation data on $L$.
Note that the obstruction to existence of a secondary grading is a class in $L,B\mathbb{Z}]=H^{1}(L,\mathbb{Z})$, and the space of secondary gradings is a torsor for $\operatorname{Map}(L,\mathbb{Z})$, hence in particular, is discrete. Thus we simply ask whether secondary gradings are equal, rather than discuss homotopies between them.
@@
Let $V$ be a complex symplectic manifold. Equip the underlying real symplectic manifold with the canonical grading constructed in Lemma 3.2. Let $L\subset V$ be a smooth complex Lagrangian. Then there is a canonical choice for the secondary grading of $L\subset V$.
@@
Let $L,V$ be as in Lemma 3.12. Fix the canonical grading/orientation data on $V$, and the canonical grading on $L$. Then secondary orientation data for $L\subset V$ is equivalent to a spin structure on $L$.